# Weakly Supervised Multiple Instance Learning for Whale Call Detection and Localization in Long-Duration Passive Acoustic Monitoring

### v1

@@ Abstract
Marine ecosystem monitoring via Passive Acoustic Monitoring (PAM) generates vast data, but deep learning often requires precise annotations and short segments. We introduce DSMIL-LocNet, a Multiple Instance Learning framework for whale call detection and localization using only bag-level labels. Our dual-stream model processes 2-30 minute audio segments, leveraging spectral and temporal features with attention-based instance selection. Tests on Antarctic whale data show longer contexts improve classification (F1: 0.8-0.9) while medium instances ensure localization precision (0.65-0.70). This suggests MIL can enhance scalable marine monitoring. Code: https://github.com/Ragib-Amin-Nihal/DSMIL-Loc
@@ I Introduction
Traditional methods rely on signal processing techniques such as spectrogram correlation [2] and energy detection [3], which compare templates with input spectrograms or sum energy within specific frequency bands. While effective for stereotyped calls in low-noise conditions, these approaches require extensive parameter tuning, are noise-sensitive, and lack mechanisms for long-duration data or weak labels. Deep learning methods [4, 5] learn features directly from spectrograms but typically focus on short audio segments ($\leq 15s$) due to computational constraints and the reliance on supervised learning with precise temporal annotations. This approach limits to capture long-range dependencies in whale calls, as highlighted by [6], which emphasized the challenges of processing continuous recordings where such calls are infrequent.
Perhaps the most significant barrier to advancing automated analysis is the reliance on detailed temporal annotations. Creating large-scale annotated datasets requires substantial expert knowledge and time investment [7]. This annotation bottleneck becomes problematic when dealing with long-duration recordings, where manually marking the precise timing of each whale call is impractical. The non-stationary nature of underwater acoustic environments, with fluctuating noise levels and varying propagation conditions over extended periods, also poses difficulties for models.
@@
This work is guided by the following research question:
@@ II-A Problem Definition
Our goal is to learn a model parameterized by $\boldsymbol{\theta}$ that maximizes $\mathbb{P}(y_{i}|\mathcal{B}_{i};\boldsymbol{\theta})$ subject to $\alpha_{i}^{j}\approx\mathbb{P}(y_{i}^{j}=1|\mathbf{x}_{i}^{j};\boldsymbol{\theta})$, where $\alpha_{i}^{j}$ are attention weights corresponding to latent instance labels. Each instance $\mathbf{x}_{i}^{j}$ is transformed into a $d$ dimesion feature vector $\boldsymbol{\phi}_{i}^{j}\in\mathbb{R}^{d}$ via:
@@
Precision $=\frac{\sum_{i}\text{card}\{t_{d}\in\mathcal{D}_{i}\text{ such that }\exists g_{k}\in\mathcal{G}_{i}\text{ with }\lvert t_{d}-g_{k}\rvert\leq\tau\}}{\sum_{i}\text{card}\{\mathcal{D}_{i}\}}$
Recall $=\frac{\sum_{i}\text{card}\{g_{k}\in\mathcal{G}_{i}\text{ such that }\exists t_{d}\in\mathcal{D}_{i}\text{ with }\lvert t_{d}-g_{k}\rvert\leq\tau\}}{\sum_{i}\text{card}\{\mathcal{G}_{i}\}}$
@@ II-C Network Architecture
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V}.\vspace*{-0.2cm}$
@@
where the softmax operation normalizes attention weights across all instances within bag $i$. During inference, the model produces both bag-level predictions: $g_{\boldsymbol{\theta}}(\mathcal{B}_{i})=\sigma\left(\sum_{j}\alpha_{i}^{j}\boldsymbol{\phi}_{i}^{j}\right)$ and instance-level localization through $\alpha^{j}_{i}$, providing a framework for whale call detection and temporal localization.
@@ II-D Loss Function
Our loss function $\mathcal{L}(\boldsymbol{\theta})$ comprises four components addressing class imbalance, temporal coherence, sparsity, and instance consistency. The primary classification loss uses focal binary cross-entropy:
$\mathcal{L}_{\text{BCE}}(y_{i},\hat{y}_{i})=-y_{i}(1-\hat{y}_{i})^{\gamma}\log(\hat{y}_{i})-(1-y_{i})\hat{y}_{i}^{\gamma}\log(1-\hat{y}_{i}).$ (1)
@@
$\mathcal{L}_{\text{temp}}=\frac{1}{M_{i}}\sum_{i}\sum_{j}(\alpha_{i}^{j}-\alpha_{i}^{j-1})^{2}.$
@@
$\mathcal{L}_{{\text{sparsity}}}=\frac{1}{M_{i}}\sum_{i}\sum_{j}|\alpha_{i}^{j}|,\mathcal{L}_{{\text{inst}}}=\frac{1}{M_{i}^{2}}\sum_{i}\sum_{j,k}\|\boldsymbol{\phi}_{i}^{j}-\boldsymbol{\phi}_{i}^{k}\|_{2}^{2}.$
@@
$\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{{\text{BCE}}}+\lambda_{t}\mathcal{L}_{{\text{temp}}}+\lambda_{s}\mathcal{L}_{{\text{sparsity}}}+\lambda_{i}\mathcal{L}_{{\text{inst}}}.$ (2)
@@ III-A Dataset
We utilize the AcousticTrends BlueFinLibrary [14], a public dataset of annotated underwater recordings from the Southern Ocean. This dataset comprises 1880.25 hours of audio across seven Antarctic sites and 11 site-year combinations, with annotations indicating the presence of Antarctic blue and fin whale vocalizations. For this study, we framed the task as binary classification (presence/absence of any whale call). The data is highly imbalanced, with a majority (73-99.9%) of recordings containing only background noise. We employed blocked cross-validation, where each site-year combination served as an independent test set in turn, ensuring robust generalization assessment to unseen recording locations and conditions. The remaining site-years were used for training and validation.
@@ III-C1 Baseline Comparisons
Model Configuration Loss Function F1-Classification Localization Precision
@@
III-C2 Classification Performance
III-C3 Localization Accuracy
Figure 3 reveals localization follows more complex patterns than classification. Highest precision (0.82) is achieved with 30s instances at 60s bag duration, while 15s instances maintain consistent precision (0.65-0.70) across different bag durations. Variable $R^{2}$ values (0.080-0.995) indicate less predictable instance-duration relationships, with wider score distributions for longer bags suggesting increased variability. The correlation matrix shows weaker and sometimes negative correlations between different bag durations (e.g., -0.22 between 300s and 1800s), highlighting cross-temporal consistency challenges.
@@ III-C4 Performance Trade-offs
    Instance Duration Effects: The trend analysis slopes show opposing patterns for classification (consistently negative slopes) versus localization (variable slopes), indicating that instance duration impacts these tasks differently.
@@
    Correlation Structure: The classification correlation matrix shows uniformly high correlations $(\geq 0.97)$ for shorter bag durations, suggesting consistent behavior. In contrast, the localization correlation matrix shows more variable relationships, with some negative correlations, indicating less predictable performance patterns.
@@ III-D Comparison with Existing Methods
To the best of our knowledge, DSMIL-LocNet represents the first MIL application to whale call detection and localization in PAM data. Recent approaches have made significant progress in automated whale call detection, but differ fundamentally in their requirements and capabilities. Table II provides a comparative overview of current state-of-the-art methods. While direct performance comparisons are challenging, our high F1 scores and precision demonstrate effectiveness in marine acoustic monitoring, providing both detection and localization within the same framework.
@@
(Ours) MIL + Attn 15s inst. + Temporal (weak) + loc.
@@ IV Conclusion
This study investigated whether Multiple Instance Learning could provide an effective framework for analyzing long-duration underwater acoustic recordings while reducing the burden of detailed temporal annotations. Our results demonstrate that incorporating longer temporal contexts through a bag-based approach significantly improves whale call detection accuracy compared to traditional fixed-window methods. DSMIL-LocNet balances the competing demands of broad temporal analysis and precise localization, with longer bag durations improving classification accuracy while maintaining temporal localization through attention-based instance selection. The dual-stream architecture effectively handles challenging marine noise conditions by combining spectral and temporal features. These findings suggest that weak supervision through MIL could transform marine acoustic monitoring, enabling more widespread deployment of automated systems while maintaining necessary precision. While our evaluation focused on whales, the approach can be applied to other species with sparsely structured acoustic signals. Future work should address the challenge of handling diverse acoustic events in complex marine soundscapes, potentially through specialized detectors or multi-class MIL approaches. This work represents a step toward scalable automated monitoring systems for marine conservation and research across multiple species and acoustic event types.

### v2

# Weakly Supervised Multiple Instance Learning for Whale Call Detection and Temporal Localization in Long-Duration Passive Acoustic Monitoring

@@ Abstract
Marine ecosystem monitoring via Passive Acoustic Monitoring (PAM) generates vast data, but deep learning often requires precise annotations and short segments. We introduce DSMIL-LocNet, a Multiple Instance Learning framework for whale call detection and temporal localization using only bag-level labels. Our dual-stream model processes 2-30 minute audio segments, leveraging spectral and temporal features with attention-based instance selection. Tests on Antarctic whale data show longer contexts improve classification (F1: 0.8-0.9) while medium instances ensure localization precision (0.65-0.70). This suggests MIL can enhance scalable marine monitoring. Code: https://github.com/Ragib-Amin-Nihal/DSMIL-Loc
@@ I Introduction
Traditional methods rely on signal processing techniques such as spectrogram correlation [2] and energy detection [3], which compare templates with input spectrograms or sum energy within specific frequency bands. While effective for stereotyped calls in low-noise conditions, these approaches require extensive parameter tuning, are noise-sensitive, and lack mechanisms for long-duration data or weak labels.
Perhaps the most significant barrier to advancing automated analysis is the reliance on detailed temporal annotations. Creating large-scale annotated datasets requires substantial expert knowledge and time investment [4]. This annotation bottleneck becomes particularly problematic for deep learning methods [5, 6], which learn features directly from spectrograms but typically focus on short audio segments ($\leq 15s$) due to both computational constraints and this requirement for precise temporal labels. The limitation to short segments restricts the ability to capture long-range dependencies in whale calls, as highlighted by [7], which emphasized the challenges of processing continuous recordings where such calls are infrequent. Furthermore, the non-stationary nature of underwater acoustic environments, with fluctuating noise levels and varying propagation conditions over extended periods, poses additional difficulties for these models.
@@
In this paper, we address two fundamental tasks in marine bioacoustic analysis: (1) call detection - determining whether whale vocalizations are present within a recording segment, and (2) temporal localization - identifying the timestamps at which these calls occur. We use the term “temporal localization” to refer specifically to determining when acoustic events happen in time, as distinct from “spatial localization” which traditionally refers to determining the physical location or direction of sound sources in underwater acoustics. “Temporal localization” commonly describes the identification of when events occur with machine learning literature in audio and video analysis. Through our MIL approach, we aim to accomplish both detection and temporal localization using only weakly labeled data, without requiring precise temporal annotations during training. This work is guided by the following research question:
@@ II-A Problem Definition
Our goal is to learn a model parameterized by $\boldsymbol{\theta}$ that maximizes $\mathbb{P}(y_{i}|\mathcal{B}_{i};\boldsymbol{\theta})$ subject to $\alpha_{i}^{j}\approx\mathbb{P}(y_{i}^{j}=1|\mathbf{x}_{i}^{j};\boldsymbol{\theta})$, where $\alpha_{i}^{j}$ are attention weights corresponding to latent instance labels for temporal localization. Each instance $\mathbf{x}_{i}^{j}$ is transformed into a $d$ dimesion feature vector $\boldsymbol{\phi}_{i}^{j}\in\mathbb{R}^{d}$ via:
@@
Precision $=\frac{\sum_{i}\text{card}\{t_{d}\in\mathcal{D}_{i}\text{ such that }\exists g_{k}\in\mathcal{G}_{i}\text{ with }\lvert t_{d}-g_{k}\rvert\leq\tau\}}{\sum_{i}\text{card}\{\mathcal{D}_{i}\}}.$
Recall $=\frac{\sum_{i}\text{card}\{g_{k}\in\mathcal{G}_{i}\text{ such that }\exists t_{d}\in\mathcal{D}_{i}\text{ with }\lvert t_{d}-g_{k}\rvert\leq\tau\}}{\sum_{i}\text{card}\{\mathcal{G}_{i}\}}.$
@@ II-C Network Architecture
$\vspace*{-0.2cm}\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V}.$
@@
where the softmax operation normalizes attention weights across all instances within bag $i$. During inference, the model produces both bag-level predictions: $g_{\boldsymbol{\theta}}(\mathcal{B}_{i})=\sigma\left(\sum_{j}\alpha_{i}^{j}\boldsymbol{\phi}_{i}^{j}\right)$ and instance-level temporal localization through $\alpha^{j}_{i}$, providing a framework for whale call detection and temporal localization.
@@ II-D Loss Function
Our loss function $\mathcal{L}(\boldsymbol{\theta})$ comprises four components addressing class imbalance, temporal coherence, sparsity, and instance consistency. The loss uses focal binary cross-entropy:
$\vspace*{-0.2cm}\mathcal{L}_{\text{BCE}}(y_{i},\hat{y}_{i})=-y_{i}(1-\hat{y}_{i})^{\gamma}\log(\hat{y}_{i})-(1-y_{i})\hat{y}_{i}^{\gamma}\log(1-\hat{y}_{i}).$ (1)
@@
$\mathcal{L}_{\text{temp}}=\frac{1}{M_{i}}\sum_{i}\sum_{j}(\alpha_{i}^{j}-\alpha_{i}^{j-1})^{2}.\vspace*{-0.15cm}$
@@
$\vspace*{-0.1cm}\mathcal{L}_{{\text{sparsity}}}=\frac{1}{M_{i}}\sum_{i}\sum_{j}|\alpha_{i}^{j}|,\mathcal{L}_{{\text{inst}}}=\frac{1}{M_{i}^{2}}\sum_{i}\sum_{j,k}\|\boldsymbol{\phi}_{i}^{j}-\boldsymbol{\phi}_{i}^{k}\|_{2}^{2}.\vspace*{-0.1cm}$
@@
$\vspace*{-0.2cm}\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{{\text{BCE}}}+\lambda_{t}\mathcal{L}_{{\text{temp}}}+\lambda_{s}\mathcal{L}_{{\text{sparsity}}}+\lambda_{i}\mathcal{L}_{{\text{inst}}}.$ (2)
@@ III-A Dataset
We utilize the AcousticTrends BlueFinLibrary [14], a public dataset of annotated underwater recordings from the Southern Ocean. This dataset comprises 1880.25 hours of audio across seven Antarctic sites and 11 site-year combinations, with annotations indicating the presence of Antarctic blue and fin whale vocalizations. For this study, we framed the task as binary classification (presence/absence of any whale call). The data is highly imbalanced, with a majority (73-99.9%) of recordings containing only background noise. We employed blocked cross-validation, where each site-year combination served as an independent test set in turn, ensuring assessment to unseen recording locations and conditions. The remaining site-years were used for training and validation.
@@ III-C2 Classification Performance
Model Configuration Loss Function F1-Classification Temp.Localization Precision
@@
III-C3 Temporal Localization Accuracy
Figure 3 reveals temporal localization follows more complex patterns than classification. Highest precision (0.82) is achieved with 30s instances at 60s bag duration, while 15s instances maintain consistent precision (0.65-0.70) across different bag durations. Variable $R^{2}$ values (0.080-0.995) indicate less predictable instance-duration relationships, with wider score distributions for longer bags suggesting increased variability. The correlation matrix shows weaker and sometimes negative correlations between different bag durations (e.g., -0.22 between 300s and 1800s), highlighting cross-temporal consistency challenges.
@@ III-C4 Performance Trade-offs
    Instance Duration Effects: The trend analysis slopes show opposing patterns for classification (consistently negative slopes) versus temporal localization (variable slopes), indicating that instance duration impacts these tasks differently.
@@
    Correlation Structure: The classification correlation matrix shows uniformly high correlations $(\geq 0.97)$ for shorter bag durations, suggesting consistent behavior. In contrast, the temporal localization correlation matrix shows more variable relationships, with some negative correlations, indicating less predictable performance patterns.
@@ III-D Comparison with Existing Methods
To the best of our knowledge, DSMIL-LocNet represents the first MIL application to whale call detection and temporal localization in PAM data. Recent approaches have made significant progress in automated whale call detection, but differ fundamentally in their requirements and capabilities. Table II provides a comparative overview of current state-of-the-art methods. While direct performance comparisons are challenging, our high F1 scores and precision demonstrate effectiveness in marine acoustic monitoring, providing both detection and temporal localization within the same framework.
@@
(Ours) MIL + Attn 15s inst. + Temporal (weak) + temp.loc.
@@ IV Conclusion
This study investigated whether Multiple Instance Learning could provide an effective framework for analyzing long-duration underwater acoustic recordings while reducing the burden of detailed temporal annotations. Our results demonstrate that incorporating longer temporal contexts through a bag-based approach significantly improves whale call detection accuracy compared to traditional fixed-window methods. DSMIL-LocNet balances the competing demands of broad temporal analysis and precise temporal localization, with longer bag durations improving classification accuracy while maintaining temporal localization through attention-based instance selection. The dual-stream architecture effectively handles challenging marine noise conditions by combining spectral and temporal features. These findings suggest that weak supervision through MIL could transform marine acoustic monitoring, enabling more widespread deployment of automated systems while maintaining necessary precision. While our evaluation focused on whales, the approach can be applied to other species with sparsely structured acoustic signals. Future work should address the challenge of handling diverse acoustic events in complex marine soundscapes, potentially through specialized detectors or multi-class MIL approaches. This work represents a step toward scalable automated monitoring systems for marine conservation and research across multiple species and acoustic event types.